# XtraLight-MedMamba for Classification of Neoplastic Tubular Adenomas

### v3

@@ 4.1 Model Parameters
Table 1 summarizes the total number of trainable parameters across Transformer-based, Mamba-based, and XtraLight mamba-based architectures. Among Transformer models, the Vision Transformer (ViT) exhibited the highest parameter count at approximately 7.4 million, followed by the Swin Transformer with 598 thousand parameters, with their hierarchical architectural structures. In contrast, Mamba-based architectures demonstrated substantially lower complexity, with parameter counts ranging from 49 thousand to 59 thousand for linear classifier variants. Further reductions were observed in the proposed XtraLight Mamba variants, where the parameter count was reduced by nearly 40–50% compared to the baseline Mamba models. The Conv. based XtraLight Mamba model had 28,329 parameters, which was the lowest count observed. Our ConvNeXt-based model, XtraLight-MedMamba, had the second-lowest count of the models studied, at only 32,073 parameters. This underscores the effectiveness of our proposed lightweight design in minimizing computational cost while maintaining strong representational performance.
4.2 Dataset
(a) (a)
(b) (b)
Figure 5: Sample images of the dataset used: (a) H&E-stained WSI tiles from the case group consisting of tubular adenomas with low-grade dysplasia from patients who subsequently develeped CRC (b) H&E-stained WSI tiles from the control group consisting of tubular adenomas with low-grade dysplasia from patients without subsequent development of CRC.
Figure 6: Control Group: TA images from patients with no CRC development during follow-up. Case Group: TA images from patients who later developed CRC.
The original whole slide images (WSIs) were tiled at $1024\times 1024$ pixel resolution with 3 color channels. They were then resized into smaller tiles of $224\times 224$ pixels with 3 color channels for model input as shown in Fig. 5. During preprocessing, a region-of-interest (ROI) filter was applied to determine whether each tile should be retained or discarded. All automatically generated ROIs were subsequently subjected to visual inspection to verify annotation accuracy. Tiles exhibiting quality issues—such as tissue folding, edge artifacts, or poor scan resolution—were excluded through manual review of WSI patch location maps. After this curation process, a total of 135,049 high-quality tiles were retained in each class. The final dataset was divided into training (70%), validation (15%), and testing (15%) subsets.
4.2.1 Data Management Strategies
Patients without known high-risk clinical factors for CRC, in which low-grade tubular adenomas were identified during screening colonoscopy, were included in the study. A total of 81 patients (41 male, 40 female), ranging in age from 54-95 years (average 70), underwent at least one screening colonoscopy with associated biopsies demonstrating tubular adenomas with low-grade dysplasia; no biopsies showed any histologic features that were indicative of high-risk progression to CRC. Patients were stratified into two cohorts: a precancer group and a control group. The case group consisted of individuals who subsequently developed CRC following screening colonoscopies in which low-grade tubular adenomas were identified. The control group comprised of individuals with no history of CRC despite having low-grade adenomas detected on one or more screening procedures. Compared to the case group, the control group had a greater average number of biopsies and a longer mean screening interval. On average, patients in the case group were 6.86 years older than those in the control group. Histologic slides from both groups containing low-grade tubular adenomas were digitized using the same Leica Aperio AT2 whole slide scanner to generate image data for this study. Fig. 6 displays the pictorial details of the two cohorts: control group and case group.
4.3 Results
Models Accuracy F1 Precision Recall
Transformer-Based Models
Vision Transformer 89.84% 0.8920 0.9519 0.8392
Swin Transformer 89.52% 0.8878 0.9548 0.8296
Mamba-Based Models with Linear Classifier
Conv. based 94.24% 0.9431 0.9324 0.9540
MBConv. based 92.44% 0.9221 0.9512 0.8948
Fused MBConv. based 93.67% 0.9388 0.9076 0.9703
ConvNext based 96.50% 0.9651 0.9606 0.9697
XtraLight Mamba Models
Conv. based 93.64% 0.9367 0.9321 0.9414
MBConv. based 93.96% 0.9117 0.9029 0.8913
Fused MBConv. based 94.24% 0.9416 0.9547 0.9289
ConvNext based (ours, XtraLight-MedMamba) 97.18% 0.9767 0.9666 0.9717
Table 2 summarizes the quantitative performance of Transformer-based architectures, Mamba-based variants, and XtraLight Mamba-based variants, employing both linear and fixed non-negative orthogonal classifiers. Among the Transformer models, which primarily utilize self-attention mechanisms to capture long-range dependencies across image patches, the Vision Transformer achieved an overall accuracy of 89.84% (F1 = 0.8920, precision = 0.9519, recall = 0.8392) with 7.39M parameters, while the Swin Transformer attained a comparable accuracy of 89.52% (F1 = 0.8878, precision = 0.9548, recall = 0.8296) using a more compact 598K parameters.
In comparison, the Mamba-based models built upon State Space Models (SSMs) that process sequential representations bidirectionally, demonstrated consistently higher performance with substantially fewer parameters. Within the linear classifier group, the ConvNeXt-based Mamba variant achieved the highest accuracy of 96.50% (F1 = 0.9651, precision = 0.9606, recall = 0.9697), outperforming the convolutional, MBConv, and Fused MBConv variants. The incorporation of the SCAB module further enhanced feature propagation, which proved beneficial for the image classification task. As the SSM effectively models hidden states over time, it excels at capturing both long- and short-range dependencies within spatial representations.
Similarly, the proposed XtraLight Mamba models maintained strong performance while significantly reducing model complexity. Our model, ConvNeXt-based XtraLight-MedMamba, achieved the best overall results, reaching an accuracy of 97.18% (F1 = 0.9767, precision = 0.9666, recall = 0.9717) with only 32K parameters, surpassing all other tested configurations. These findings highlight that integrating ConvNeXt-style inverted bottlenecks within the Mamba framework provides an optimal balance between efficiency and accuracy, outperforming both Transformer and other Mamba variants.
Table 3 presents the normalized confusion matrix for the proposed XtraLight-MedMamba model. The classifier achieved exceptionally high discriminative performance, correctly identifying 97.7% of positive samples (true positives) and 96.7% of negative samples (true negatives). Only a small fraction of the cases were misclassified, with 2.3% false negatives and 3.3% false positives. These results demonstrate the strong sensitivity and specificity of the model, confirming its robustness and reliability in distinguishing between positive and negative classes. The high diagonal dominance of the matrix further reflects the effectiveness of the learned feature representations and the generalizability of the model throughout the dataset.
Table 3: Confusion matrix for XtraLight-MedMamba
Predicted Positive Predicted Negative
[…]
4.4 Model Interpretability and Visual Explanation
(a) (a)
(b) (b)
To aid interpretation of the decision-making behavior of the proposed XtraLight-MedMamba model, Gradient-weighted Class Activation Mapping (Grad-CAM)  [20] was used to visualize the spatial regions contributing most strongly to classification. As illustrated in Fig. 7, the generated activation maps highlight the discriminative tissue areas that influenced the model’s predictions. Regions with warmer colors (red–yellow) correspond to areas of greater interest, whereas cooler colors (blue–purple) indicate lesser contribution. The Grad-CAM overlays demonstrate that the network predominantly focuses on histologically relevant structures—such as atypical epithelial glands, nuclear crowding, and stromal interfaces—rather than background artifacts. This highlights how model activations and pathologically significant features support the interpretability and reliability of the model, confirming that XtraLight-MedMamba learns clinically meaningful representations rather than false correlations.
A closer look at the Grad-CAM visualizations in Fig. 7 shows that the model’s attention is drawn mainly to nuclear architecture. In both case and control images, the highlighted regions consistently align with areas rich in nuclear material, especially where the nuclei appear pseudostratified or show subtle irregularities in their basal orientation. In case samples, these activations often correspond to regions with more pronounced pseudostratification and occasional loss or disorder of polarity, while control samples display well-aligned nuclei with layering to case nuclei. Together, these observations suggest that XtraLight-MedMamba differentiates tissues primarily by recognizing patterns of nuclear arrangement and polarity—the same microscopic cues pathologists use to identify epithelial atypia and assess disease progression.
@@ 4.5 Mismatched predicted outputs of XtraLight-MedMamba
Figure 8: Representative mismatched predictions made by XtraLight-MedMamba. (a) Target: Case, Predicted: Control. The glands in these sections show preserved architecture with evenly spaced crypts and relatively consistent nuclear alignment. Cytologic atypia in these sections are subtle, showing mild nuclear enlargement and crowding, morphological features closely resembling low-risk tubular adenomas. The lack of overt glandular complexity or pronounced dysplasia likely contributed to the model misclassifying the sample as Control. (b) Target: Control, Predicted: Case. In contrast, these Control samples exhibit focal areas of gland crowding, mild loss of polarity, and mild pseudostratification of the nuclei with darker nuclei. These morphological patterns tend to be more commonly associated with higher-risk lesions, which may have led the model to weigh these as Case associated features.
In Fig. 8 (a), the patch clearly demonstrates morphological features of tubular adenoma: nuclear pseudostratification, hyperchromatic elongated nuclei, and goblet cell depletion. Despite the morphologic hallmarks being present, the model misclassified this region as class control. This could reflect the model’s limited sensitivity to focal dysplasia, particularly in areas with mixed histology, where adjacent non-dysplastic crypts can mask subtle neoplastic changes. These results highlight the need for enhanced model training with finer-grained annotations and increased representation of early or borderline dysplasia cases.
In Fig. 8 (b), the region is histologically consistent with slight adenomatous changes from the control group and was misclassified by the model as a tubular adenoma that progressed to CRC (case group). The crypts are well-formed, with preserved spacing, abundant goblet cells, and basally aligned nuclei, lacking pseudostratification, along with no cytologic atypia indicative of a high-grade dysplasia. Morphologically, there are no other signs of progression to CRC. This false positive might be due to subtle visual cues such as epithelial overcrowding near the tissue edge or darker nuclear staining, which the model may overfit during training.

### v4

@@ 4 Experimental Setup and Results
4.1 Dataset Curation
(a) (a)
(b) (b)
Figure 5: Sample images of the dataset used: (a) H&E-stained WSI tiles from the case group consisting of tubular adenomas with low-grade dysplasia from patients who subsequently developed CRC (b) H&E-stained WSI tiles from the control group consisting of tubular adenomas with low-grade dysplasia from patients without subsequent development of CRC.
Figure 6: Illustration of post-colonoscopy outcomes. Control Group: TA images from patients with no CRC development during follow-up. Case Group: TA images from patients who later developed CRC.
A subset of WSIs was manually annotated by expert pathologists to mark Regions of Interest (ROIs) containing adenomatous epithelium. Three-color channel tiles of $1024\times 1024$ pixels containing tissue from the WSI were extracted. Each tile was labeled ROI-positive if it overlapped an annotated region, and ROI-negative otherwise. The tiles were resized to $224\times 224\times 3$ to meet the input requirements of the EfficientNetV2S [42]. The model was trained on annotated WSI tiles to predict whether unseen tiles contained relevant tissue for this study.
Following training, the model was applied to the full set of WSIs. These WSIs were tiled at $1024\times 1024\times 3$ and then resized into smaller tiles of $224\times 224\times 3$ pixels for model input as shown in Fig. 5. During preprocessing, the trained EfficientNetV2S model evaluated each tile to determine whether it should be retained or discarded. The automatically generated ROIs were visually assessed for annotation accuracy. Tiles affected by quality issues, including tissue folding, edge artifacts, or suboptimal scan resolution, were excluded based on manual examination of the WSI patch location maps. Following curation, a total of 135,049 high-quality tiles were retained per class. The resulting dataset was subsequently split into training (70%), validation (15%), and testing (15%) subsets [31].
4.1.1 Data Management Strategies
Patients with low-grade TAs detected at screening colonoscopy were selected for the study after excluding individuals with known high-risk CRC predisposition or histologic features associated with advanced neoplasia. The cohort included 81 patients, comprising 41 males and 40 females, ranging in age from 54 to 95 years (average 70 years) whose biopsies showed TAs with low-grade dysplasia only and no histologic features suggestive of high-risk progression to CRC. The patients were grouped into case and control cohorts based on longitudinal outcomes. The case cohort included patients who subsequently developed CRC after screening colonoscopies in which low-grade TAs were identified, whereas the control cohort included patients without CRC development during follow-up despite detection of low-grade TAs on one or more examinations. The control group had more biopsies and longer mean surveillance intervals, and patients from the case group were on average 6.86 years older than those in the control group. Fig. 6 shows the pictorial details of the two cohorts: the control and case groups [31]. For further details regarding the dataset, including patient inclusion/exclusion criteria and cohort definition, please refer to [31] describing the study design.
4.1.2 Data Acquisition
Histological slides from both groups containing low-grade tubular adenomas were digitized using the same Leica Aperio AT2 [18] instrument, an automated whole-slide scanner capable of unattended scanning of up to 400 standard 75mm x 25mm glass slides through an integrated AutoLoader to generate image data for this study. The spatial resolution of the WSIs was 50,000 pixels per inch and $0.25\mu m/pixel$ at 40x. A Plan-Apochromat 20x/0.75 NA objective lens was used with an integrated 2x optical magnification changer, enabling 40x scanning with a single objective lens. Scanning was performed using a line-scan imaging method with fully automated tissue detection and automatic focus. The stored WSIs are 24-bit color images at gigapixel resolution, and the slides were generated in a contiguous pyramidal TIFF format, which was viewable through Aperio ImageScope software.
4.2 Training Method
All models, including transformer-based models, and mamba-based models with linear classifier and XtraLight mamba models were trained for 100 epochs under a common baseline configuration using Stochastic Gradient Descent (SGD) with a momentum of 0.99, learning rate of $1\times 10^{-5}$ and batch size of 256. The loss function was binary cross-entropy. Salt and pepper noise was randomly injected into the latent feature space during training rather than being directly applied to the input images to improve robustness and mitigate overfitting. The One-Cycle Learning Rate (OneCycleLR) scheduler, along with Stochastic Weight Averaging (SWA) to stabilize training, was incorporated to improve optimization stability. All experiments were implemented in PyTorch and executed on a NVIDIA GeForce RTX Titan GPU.
4.3 Results
Models Accuracy F1 Precision Recall
Transformer-Based Models
Vision Transformer 89.84% 0.8920 0.9519 0.8392
Swin Transformer 89.52% 0.8878 0.9548 0.8296
Mamba-Based Models with
Linear Classifier
Conv. based 94.24% 0.9431 0.9324 0.9540
MBConv. based 92.44% 0.9221 0.9512 0.8948
Fused MBConv. based 93.67% 0.9388 0.9076 0.9703
ConvNeXt based 96.50% 0.9651 0.9606 0.9697
XtraLight Mamba Models
Conv. based 93.64% 0.9367 0.9321 0.9414
MBConv. based 93.96% 0.9117 0.9029 0.8913
Fused MBConv. based 94.24% 0.9416 0.9547 0.9289
ConvNeXt based 97.18% 0.9767 0.9666 0.9717
(ours, XtraLight-MedMamba)
Table 1 summarizes the quantitative performance of Transformer-based architectures, Mamba-based variants, and XtraLight Mamba-based variants, employing both linear and fixed non-negative orthogonal classifiers. Among the Transformer models, which primarily utilize self-attention mechanisms to capture long-range dependencies across image patches, the Vision Transformer model achieved an overall accuracy of 89.84% (F1 = 0.8920, precision = 0.9519, recall = 0.8392) with 7.39M parameters, while the Swin Transformer attained a comparable accuracy of 89.52% (F1 = 0.8878, precision = 0.9548, recall = 0.8296) using a more compact 598K parameters.
In comparison, the Mamba-based models built upon State Space Models (SSMs) that process sequential representations bidirectionally, demonstrated consistently higher performance with substantially fewer parameters. Within the linear classifier group, the ConvNeXt-based Mamba variant achieved the highest accuracy of 96.50% (F1 = 0.9651, precision = 0.9606, recall = 0.9697), outperforming the convolutional, MBConv, and Fused MBConv variants. The incorporation of the SCAB module further enhanced feature propagation, which proved beneficial for image classification. As the SSM effectively models hidden states over time, it excels at capturing both long- and short-range dependencies within spatial representations.
Similarly, the proposed XtraLight Mamba models maintained strong performance while significantly reducing model complexity. Our model, ConvNeXt-based XtraLight-MedMamba, achieved the best overall results, reaching an accuracy of 97.18% (F1 = 0.9767, precision = 0.9666, recall = 0.9717) with only 32K parameters, surpassing all other tested configurations. These findings highlight that integrating ConvNeXt-style inverted bottlenecks into the Mamba framework strikes an optimal balance between efficiency and accuracy, outperforming both the Transformer and other Mamba variants.
Table 2 presents the normalized confusion matrix for the proposed XtraLight-MedMamba model. The classifier achieved exceptionally high discriminative performance, correctly identifying 97.7% of positive samples (true positives) and 96.7% of negative samples (true negatives). Only a small fraction of the cases were misclassified, with 2.3% false negatives and 3.3% false positives. These results demonstrate the model’s strong sensitivity and specificity, confirming its robustness and reliability in distinguishing between positive and negative classes. Strong diagonal dominance confirms effective feature learning and generalization.
Table 2: Confusion matrix for XtraLight-MedMamba
Predicted Positive Predicted Negative
[…]
4.4 Model Interpretability and Visual Explanation
(a) (a)
(b) (b)
To aid interpretation of the decision-making behavior of the proposed XtraLight-MedMamba model, Gradient-weighted Class Activation Mapping (Grad-CAM)  [35] was used to visualize the spatial regions that contribute most to classification. As illustrated in Fig. 7, the generated activation maps highlight the discriminative tissue areas that influenced the model’s predictions. Regions with warmer colors (red–yellow) correspond to areas of greater interest, whereas cooler colors (blue–purple) indicate lesser contribution. The Grad-CAM overlays demonstrate that the network predominantly focuses on histologically relevant structures—such as atypical epithelial glands, nuclear crowding, and stromal interfaces, rather than background artifacts. This alignment between activation patterns and pathologically relevant structures supports interpretability and reliability, confirming that XtraLight-MedMamba learns clinically meaningful representations rather than false correlations.
A closer look at the Grad-CAM visualizations in Fig. 7 shows that the model’s attention is drawn mainly to nuclear architecture. In both case and control images, the highlighted regions consistently align with areas rich in nuclear material, especially where the nuclei appear pseudostratified or show subtle irregularities in their basal orientation. In case samples, these activations often correspond to regions with more pronounced pseudostratification and occasional loss of polarity, whereas control samples display well-aligned nuclei with layering comparable to case nuclei. Together, these observations suggest that XtraLight-MedMamba differentiates tissues primarily by recognizing patterns of nuclear arrangement and polarity the same microscopic cues pathologists use to identify epithelial atypia and assess disease progression.
@@ 4.5 Model Parameters
Table 3 summarizes the total number of trainable parameters across Transformer-based, Mamba-based, and XtraLight Mamba-based architectures. Among Transformer models, the Vision Transformer (ViT) had the highest parameter count, at approximately 7.4 million, followed by the Swin Transformer with 598 thousand parameters, owing to their hierarchical architectures. In contrast, Mamba-based architectures demonstrated substantially lower complexity, with parameter counts ranging from 49,000 to 59,000 for linear classifier variants. Further reductions were observed in the proposed XtraLight Mamba variants, with parameter counts reduced by nearly 40–50% compared to the baseline Mamba models. The Conv.-based XtraLight Mamba model, with 28,329 parameters, had the fewest parameters among the models. Our ConvNeXt-based model, XtraLight-MedMamba, has the second-lowest count of the models studied, at only 32,073 parameters. This underscores the effectiveness of our proposed XtraLight design in minimizing computational cost while maintaining strong representational performance.
4.6 Ablation Studies and Discussion
F-1 Score and Recall were used for the ablation studies due to their robustness to class imbalance and clinical relevance. In this study, recall, also known as sensitivity, quantifies the model’s ability to correctly identify high-risk adenomas, making false negatives particularly harmful. F1-score mitigates the potential class imbalance by providing a balanced assessment of precision and recall. Although accuracy provides completeness, F1-score and recall offer more clinically relevant evaluation and a more informative assessment of discriminative performance when comparing architectural and optimization variants.
4.6.1 Effect of momentum on model performance
An ablation study was conducted to assess the impact of momentum values 0.8, 0.85, 0.9, 0.95, 0.99, and 1.0 on model performance. As illustrated in Fig. 8, increasing momentum progressively improved the F1-Score and Recall, with optimal performance achieved at 0.99, followed by a slight dip at 1.0. By accumulating the exponential moving average (EMA) of past gradients, momentum smooths stochastic gradient updates. This enhances convergence along stable descent directions, reduces oscillations, and improves generalization ability of the model [28], [29], [15], [41]. Although momentum at 1.0 shows descent performance, the slight dip compared to momentum at 0.99 suggests that reduced damping of gradient noise causes the optimizer to overshoot the optimal minimum at the extreme limit. Overall, high momentum effectively increases the step size in consistent directions, facilitating convergence toward flatter minima and improving generalization by escaping sharp local minima [13], [15].
Figure 8: Analysis of momentum ranging from 0.80 to 1.0 versus F1-Score and Recall, where the momentum setting of 0.99 is highlighted, demonstrating the best overall performance among the evaluated configurations.
4.6.2 Effect of learning rate on model performance
Impact of adjusting the learning rate from $1\times 10^{-7}$ to $1\times 10^{-3}$ on F1-Score and Recall is shown in Fig. 9. Improvement in F1-Score and Recall was seen as the learning rate was increased from $1\times 10^{-7}$ to $1\times 10^{-5}$, with optimal performance achieved at $1\times 10^{-5}$. At lower learning rates, performance is severely constrained, leading to slower convergence and insufficient parameter updates caused by suboptimal optimization [4]. Alternatively, increasing the learning rate beyond the optimal value can lead to overshooting of minima and unstable updates, degrading model generalization [10], [15]. The performance dip observed at $1\times 10^{-4}$ and marginal recovery at $1\times 10^{-3}$ suggest optimization instability at larger update steps, where they destabilize convergence and hinder optimization. The peak performance achieved at $1\times 10^{-5}$ learning rate represents the optimal balance between effective step size and convergence precision required for the generalization ability of the model.
Figure 9: Analysis of learning rate ranging from $1\times 10^{-7}$ to $1\times 10^{-3}$ versus F1-Score and Recall, where learning rate setting of $1\times 10^{-5}$ is highlighted, demonstrating the best overall performance among the evaluated configurations.
4.6.3 Effect of optimizers on model performance
Figure 10: Performance comparison of SGD, Adam, and AdamW optimizers evaluated using F1-Score and Recall, where the SGD optimizer demonstrated improved classification performance.
An ablation study was conducted to compare different optimizers, i.e., SGD, Adam, and AdamW, as shown in Fig. 10. Adaptive optimizers such as Adam accelerate convergence by adjusting per-parameter learning rates [17], settling into sharper local minima, causing a generalization gap [45], [15]. AdamW, which decouples weight decay from gradient updates, improves regularization [23] but did not outperform SGD based on our ablation experiments. SGD coupled with momentum often converges toward flatter minima, which are associated with improved generalization performance [13], thereby providing favorable optimization and generalization for histopathology classification.
4.6.4 Effect of different classifier configurations on model performance
An ablation study of model performance on different classifier configurations was conducted as shown in Table 4. The configuration of the proposed model, XtraLight-MedMamba, with three linear and FNO layers, provides an optimal balance between structural regularization and flexibility. The all-FNO configuration imposes constrained or fixed orthonormal classifier directions, which may restrict the ability of the model to adapt heterogeneous patterns from feature representation, potentially limiting effective margin formation and sensitivity [27], [51]. Alternatively, the all-hadamard [7], [5] configuration promotes the decorrelation of the feature and inter-class separability but lacks the structured geometric alignment introduced by FNO layers [34]. The proposed hybrid design, consisting of three linear and two FNO layers, leverages the combination of adaptive linear layers to capture specific feature geometry under the orthogonality constraints imposed by the FNO layers, yielding improved feature representation, wider decision margins, and superior generalization performance.
Table 4: Performance comparison of different classifier configurations within the proposed framework.
Classifiers Accuracy F1 Precision Recall # params
All FNO 80.95% 0.8392 0.7262 0.7938 25,884
3 Linear 2 Hadamard 87.77% 0.8884 0.8173 0.8730 32,401
All Hadamard 92.44% 0.9252 0.9163 0.9342 26,260
XtraLight-MedMamba 97.18% 0.9767 0.9666 0.9717 32,073
(ours, 3 Linear, 2 FNO)
4.6.5 Model parameters
The number of trainable parameters for each classifier configuration is shown in Table 4. Due to structural differences, the parameter sizes of the learnable linear classifier, FNOClassifier, and the Hadamard classifier differ. Configurations such as all-linear layers require more parameters due to fully trainable matrices, thereby expanding the model’s representational capacity. FNO layers exhibit minimal additional trainable parameters due to their fixed orthonormal classifier directions. This imposes structured geometric constraints and reduces model complexity. Hadamard layers rely on structured transformations that produce comparatively fewer learnable parameters while enhancing feature decorrelation. The findings demonstrate that higher parameter complexity alone is insufficient to explain performance improvements. Although the proposed model, XtraLight-MedMamba (3-linear and 2-FNOClassifier) contains the second highest number of parameters at 32,073, it achieves the superior performance compared to other classifier configurations.
@@ 4.7 Mismatched predicted outputs of XtraLight-MedMamba
Figure 11: Representative mismatched predictions made by XtraLight-MedMamba. (a) Target: Case, Predicted: Control. The glands in these sections retain their architecture, with evenly spaced crypts and relatively consistent nuclear alignment. Cytologic atypia in these sections is subtle, with mild nuclear enlargement and crowding, morphological features that closely resemble those of low-risk tubular adenomas. The lack of overt glandular complexity or pronounced dysplasia likely contributed to the model misclassifying the sample as Control. (b) Target: Control, Predicted: Case. In contrast, these Control samples exhibit focal areas of gland crowding, mild loss of polarity, and mild pseudostratification of the nuclei with darker nuclei. These morphological patterns are more commonly associated with higher-risk lesions, which may have led the model to treat them as case-associated features.
In Fig. 11 (a), a misclassified case is illustrated, showing tiles that clearly demonstrate low-grade adenomatous features, including nuclear crowding with elongation and hyperchromasia, nuclear pseudostratification, and goblet cell depletion. This misclassification likely reflects the patchy nature of dysplasia and its presence within mixed histology, where focal dysplasia and adjacent benign crypts may obscure subtle neoplastic changes and dilute the morphological features on which the model relies. This finding highlights the need for finer-grained annotations, including tighter ROIs and more examples of early dysplasia during training.
In Fig. 11 (b), the tiles belong to the control group; however, the model misclassified them as the case group that progressed to CRC despite the largely preserved architecture. Histologically, the glands are orderly and well-spaced, crypts are well-formed and well-spaced, goblet cells are plentiful, and nuclei are basally aligned without convincing pseudostratification or high-grade cytologic atypia. There are no obvious morphologic features that would suggest progression toward CRC. One possibility is that the model is being influenced by non-biological cues, such as epithelial crowding near the tissue edge or darker nuclear staining, rather than true progression-related morphology.